OpenStudy (anonymous):

The curve with equation y = 2x2 + bx + 8 has a turning point at x = 3. Find the value of b, show working please.

OpenStudy (cwrw238):

y' = 4x + b = 0 at turning points

OpenStudy (cwrw238):

when x = 3 y' = 4(3) + b = 0 12 + b = 0 b = -12